A major-league pitcher can throw a ball in excess of 41.0 m/s. If a ball is thrown horizontally at this speed, how much will it drop by the time it reaches a catcher who is 17.0 m away from the point of release? (Write on paper)
Answer #1

  Vx = 41.0 m/s (assuming a perfectly horizontal throw)

Using d = r x t
Solving for t
t = (17/41.0)
t = 0.4146 seconds

Y = Yo + Voy t - (1/2)gt^2
Substituting:
Y = 0 + 0 - (4.9)(0.4146)^2
y = 0.5422 m